OpenStudy (anonymous):

15-5(3*2-4)/10-2(4*5-16

OpenStudy (bahrom7893):

15-5(6-4)/10-2(20-16)

OpenStudy (bahrom7893):

15-5*2/10 - 2*4= 15-1-8 = 15-9 = 6

OpenStudy (anonymous):

-2003.5

OpenStudy (bahrom7893):

what's wrong with u infauzan?

OpenStudy (anonymous):

its set up as a fraction

OpenStudy (anonymous):

i got 6

OpenStudy (bahrom7893):

ohh wait did u mean: [15-5(3*2-4)]/[10-2(4*5-16)]

OpenStudy (anonymous):

ooooooooh. then 10 yes?

OpenStudy (bahrom7893):

then it is: [15-5*2]/[10-2*4]=5/2 = 2.5

OpenStudy (anonymous):

so 2.5

OpenStudy (anonymous):

im confused

OpenStudy (bahrom7893):

roseanne u need to take a picture of that problem haha this is weird... just use parenthesis all the time, extra ones help..

OpenStudy (bahrom7893):

Okay wait hold on roseanne.. ima ask u what u meant..

OpenStudy (bahrom7893):

okay almost there..

OpenStudy (anonymous):

dont't forget how to use substract, addition, multiply, and divide this is the key of he answer

OpenStudy (anonymous):

and the sign of (.....) this is the primary

OpenStudy (bahrom7893):

http://i54.tinypic.com/f9piyc.jpg Okay take a look at that picture.. which one did u mean? 1 or 2?

OpenStudy (anonymous):

so the final answer?

OpenStudy (bahrom7893):

well which one did u mean? 1 or 2? from that picture? which problem is urs?

OpenStudy (radar):

Sometimes the pursuit of an answer overplays the importance and satisfaction that comes from the development of the solution. (radar @ 2011)